A shoe company is interested in estimating the average amount customers spent at their new store in a mall. A sample of 25 customers produced a mean spending of RM150 and with standard deviation of RM20. A 95% confidence interval for the true mean is:

A. RM150 ± 164.80 C. RM150 ± 8.26 B. RM150 ± 8.24 D. RM150 ± 165.12
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = RM150

sample standard deviation = s = RM20

sample size = n = 25

Degrees of freedom = df = n - 1 = 25 - 1 = 24

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,24 = 2.064

At 95% confidence level of \mu is

\bar x\pm t\alpha/2,df * (s /\sqrtn)

150 \pm 2.064 * (20 / \sqrt 25)

RM150 \pm 8.26

option C.